OpenStudy (anonymous):

Find an equation of the line containing the given pair of points. (-2,-3) and (-6,-5) The equation of the line in slope-intercept form is y=

OpenStudy (anonymous):

first find the slope. do you know how to do that?

OpenStudy (anonymous):

Not really, algebra confuses me completely. I know about the y=mx+b and understand it if it has the m in the problem.

OpenStudy (anonymous):

you'd use point-slope formula: y2 - y1 over x2-x1 set one pair of points to (x1, y1) and the other pair to (x2,y2) and plug the values into the equation and find the slope. what is the slope of those two points?

OpenStudy (anonymous):

-13/2

OpenStudy (anonymous):

(-2, -3) (-6,-5) (x1, y1) (x2, y2) \[\frac{-5 - (-3)}{-6 - (-2)}\]

OpenStudy (anonymous):

should be \[\frac{1}{2}\]

OpenStudy (anonymous):

now use the slope-intercept equation y = mx+b to find the y-intercept (or "b" from the equation. plug in 1/2 for m and any ONE of the pair of points for x and y

OpenStudy (anonymous):

b=-2

OpenStudy (anonymous):

now use those to substitute into y = mx+b -2 for b 1/2 for m y = 1/2x -2